If the required reserve ratio is reduced from 10% in 2020 to 5% in 2021 and the central buys $29.0M worth of securities in 2021, what will be the maximum potential increase in the money supply


1
Expert's answer
2022-03-23T13:47:40-0400

money multiplier="\frac{1}{r}"

change in ratio=10-5=5%

therefore

money multiplier="\frac{1}{0.05}"

=20

then the potential maximun money supply is

20"\times"$29

=$580